Instrument Log 

BSRN / BRS C1 Lamont, OK

Date

Technician

Comments

6/10/16

JES

Radiometers are scheduled to be replaced for calibration. R/R the DD, NIP, and DS. Edited the SIRS program and loaded it to the data logger at 1918 GMT. Connected the RSU to all of the radiometer connectors and recorded the readings. Old S/N: DD - 33238, NIP - 29851, DS 30812, New S/N: DD - 33785, NIP - 34504, DS 30811. RTS @ 1945 GMT.

12/31/15

JEM

The central facility experienced an ice storm and the radiometers and suntracker are covered with ice. Removed all ice from radiometers, ventilator housings and suntracker. Cleaned the domes at 1516 GMT and RTS @1520 GMT.

9/11/15

JES

DIR is scheduled to be replaced for calibration. R/R the *DIR. Edited the program and loaded it to the data logger at 1957 GMT. The new program is called SGP_BRS_C01_F. RTS @ 2000 GMT.

*Old DIR S/N - 30357F3, New DIR S/N - 30013F3.

8/14/15

JES

*Radiometers are scheduled to be replaced for calibration. R/R the DD, NIP, and DS. Edited the BRS program and loaded it to the data logger at 1924 GMT. Confirmed that the BRS program is running correctly. The new BRS program is named SGP_BRS_C01_E. RTS @ 1930 GMT.

*Old S/N: DD - 33785, NIP - 34504E6, DS - 30811F3

*New S/N: DD - 33238, NIP - 29851E6, DS - 30812F3

6/10/15

JEM

Every time there is a power bump or outage the BRS program is restarting. Checked the backup batteries in the CR3000 and they were dead. Realized that in order for the regulator that charges these batteries to work the input voltage must be between 17 and 28VDC. The logger is currently running on a 12VDC power supply. Added a 24V power supply to the BRS so the backup batteries would charge and program restarts will be limited to extended power outages. RTS @1800 GMT.

6/10/15

JEM

In DQPR #4657 the mentor requested that the instrument alignment, levels and ventilators be checked. NIP alignment was perfect. Suntracker level was slightly off. Made adjustments to suntracker level. After leveling the tracker the DIR and DD level had to be slightly adjusted to compensate for the tracker adjustment. Checked the DS level and the bubble was within the circle but all the way to one side, so I went ahead and made an adjustment to perfectly center the bubble. Ventilators and ventilator screens were OK. RTS @1630 GMT.

5/21/15

JEM

DQOR #4657 requests the instrument cables and connections be checked regarding an issue with the derived short direct normal difference plots. Checked and reseated all connections. No problems found with instrument or cables. RTS @1555 GMT.

4/22/15

JEM

DQPR #4657 says starting on 03/07 the SGP BRS derived short direct normal difference plot shows a sudden shift and drop in derived irradiance, as well as a lower ratio. This near-inverse trend continues indefinitely. A check of the instrument revealed that the incorrect program (SGP_BRS_C01_C.CR3) was running. The proper (SGP_BRS_C01_D.CR3) was on logger but was not actively running. Removed the old program and set the correct program to "Run" and "Run on power-up" at 2100 GMT. RTS @2100 GMT.

2/10/15

JEM

SDS has been unable to collect from the BRS for 9 hours. Troubleshooting found the central cluster UPS had failed. This caused the MOXA fiber hub to lose power. R/R the UPS batteries. Powered the UPS back on and communications links were restored at 1600 GMT. RTS @1600 GMT.

7/25/14

JEM

The SIRS ventilators are being upgraded and switched from AC to DC. The *DS, **DD and ***NIP also need to be replaced for calibration. Stopped the SIRS program at 1307 GMT. Radiometers were switched out and ventilators were rewired and upgraded between 1307 and 1825 GMT. Loaded the new SIRS program @1830 GMT. Verified serial numbers cal. factors and readings. RTS @1900 GMT.

Old: *DS - 30812 **DD - 33238 ***NIP - 29851

New: *DS - 30811 **DD - 33785 ***NIP 34504

4/24/14

KT

Work order requested that the Plexiglass ring on the BRS DS be replaced. At 1535 GMT removed and replaced the the Plexiglass ring on the BRS DS. RTS at 1550 GMT.

4/18/14

JEM

The DD ventilator fan is not working. R/R DD ventilator. Leveled radiometer @1630 GMT and RTS.

2/24/14

CAW

BRS MFRSR head due 90 day calibration check. Removed MFR head at 1350 GMT and completed the 90 day calibration check. Reinstalled the MFR head at 1915 GMT.

12/23/13

JEM

The DD, DS and DIR are covered with ice. Cleared ice from the radiometers between 1800 and 1830 GMT. RTS @1830 GMT.

12/06/13

MS

Sun Tracker was not tracking at 1557 GMT. Power cycled Sun Tracker at 1612 GMT, Sun Tracker is now tracking correctly.

10/17/13

JEM

DQPR #3869 indicates that beginning at the start of the specified period (9/23/13 @1800) the short_direct_normal (NIP instrument) measurement signal occasionally drops out. Troubleshooting found a bad connection in the NIP connector. Repaired this connection @1500 GMT. Readings are OK now. RTS @1510 GMT.

10/8/13

JEM

A communications upgrade needs to be done at E13 and C1. During this upgrade power was disconnected so that power cords could be rerouted. BRS data is affected from the 1800 GMT to 1830 GMT time frame.

7/2/13

JEM

The DS, DD and NIP need to be changed out for calibration. Stopped the datalogger program @1412 GMT. R/R the *DS from 1420-1423 GMT. R/R the **DD from 1423-1426 GMT. R/R the NIP from 1426-1429 GMT. Loaded the new datalogger program @1520 GMT. Deleted the old program from the datalogger. Verified serial numbers and readings. RTS @1530 GMT.

Old - *30811 **33785 ***34504

New - *30812 **33238 ***29851

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm. Powered central cluster UPS off at 1730 GMT. Power was restored at 1920 GMT. Powered UPS on at 1945 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

10/30/12

JEM

The paint is chipping off of the suntracker arms so they need to be replaced with freshly painted ones. R/R suntracker arms between 1915 GMT and 2000 GMT. RTS @2000 GMT.

6/21/12

JES

Radiometers are scheduled to be replaced for calibration. R/R *DD, **NIP, and ***DS radiometers, edited SIRS program and loaded at 1459 GMT. RTS @ 1503 GMT.

*Old - 33238 *New - 33785 **Old - 29851E6 **New - 34504E6

***Old - 30812F3 ***New - 30811F3

5/15/12

PAD

BSRN suntracker needs alignment adjustment. Connected laptop to suntracker. Suntracker clock was 40 seconds slow. Adjusted clock and adjusted zenith +.2 deg. RTS at 1935 GMT.

5/2/12

CAW

BRS/C-1 MFRSR head due 90 calibration check. Shut down C-1 MFRSR and removed head for calibration. Completed calibration check and reinstalled head. RTS at 1825 GMT.

2/28/2012

JEM

A new program needs to be loaded to the datalogger. The previous program had incorrect serial numbers and calibration coefficients. Loaded the new program to the datalogger @1831 GMT. RTS @1840 GMT.

2/23/12

JEM

The instrument mentor has made a program change and the new program needs to be loaded. Deleted the old program from the CR3000 memory and loaded the new program @2153 GMT.

2/21/12

JEM

The BRS is schedule for a system upgrade. Powered the old system down @1935 GMT. Removed old components and installed *CR3000. Loaded the CR3000 program @2041 GMT. Readings were observed normal. Notified mentor that upgrade was complete. RTS @2045 GMT.

*Old DL - X4648 Old Multiplexer - 16194 Old PS12 - 12306

*New CR3000 - 5403

12/15/11

PAD

DQPR 2702 asks ops to check SN and calibration coefficients of DS, DD, and NIP instruments. It appears at least one of them, DD, has incorrect information since last instrument change out for calibration. DS and NIP SN and cal coefficients check out correct. Changed DD calibration factor from 114.74 to 109.51. Connected with BRS datalogger and loaded updated program. RTS at 2046 GMT.

9/13/11

PAD/KT

Radiometer ventilator fan screens need to be cleaned. Removed mounting bolts and cleaned screens on DIR, and DD. DS was checked earlier. Checked level on instruments and changed desiccant as needed. RTS at 1540 GMT.

9/1/11

PAD

DQPR 2702 states "Downwelling shortwave values appear to drop below the empirical limit (-10 W/m^2) at night, and during the day the values are less than the calculated downwelling shortwave by about 20-30 W/m^2. Could the ops please check the condition of the ventilator fan for any obstruction?" At 1510 GMT, inspected and cleaned ventilator fan screen. Secured and leveled instrument. RTS at 1525 GMT.

6/14/2011

JEM

The BRS is not transferring data to the SDS. The rocketport that the BRS is connected too has already been replaced to resolve other communication issues. Further troubleshooting found that one of the optical bit drivers has failed. R/R the bit driver. Communications are now restored. RTS @2130 GMT.

6/14/2011

JEM

A lightning strike on the central cluster on 6/12/11 has damaged several instruments. Troubleshooting found that the NIP has failed due to the lightning strike. The *DS, **DD and ***NIP are scheduled to be changed out for calibrations so that will be done now as well. Loaded the new program to the datalogger @1739 GMT. At 1740 GMT the SIRS was back to normal operation. RTS @1740 GMT.

Old - DS-30811F3 DD-33785 NIP-36316E6

New - DS-30812F3 DD-33238 NIP-29851E6

6/13/11

PAD

BRS suntracker is at home position. Power cycled suntracker but it did not start. Storms with lightning had moved across CF area at time of instrument outage. TS found blown fuse on inlet power to instrument. Replaced fuse. Connected laptop with instrument to reset configuration file. After connecting with laptop, instrument restarted. Checked and adjusted clock. RTS at 1610 GMT.

4/15/11

JEM

The instrument has not collected since the 0400 hour. Due to other instruments that are down the source of the problems seems to be the Rocketport hub. The Rocketport power light is flashing. Power cycling the Rocketport hub did not correct the issue. Notified the IT department of the issue.

4/4/11

CAM

Last successful SDS collection for BRS, 10MFR and MFRC1 was at 1800 GMT on 4/3. Power cycled the Rocket Port hub that all 3 instruments share at 1250 GMT on 4/4. The following collection was successful at 1316 GMT.

3/10/11

PAD

Suntracker is slightly out of alignment. Connected with suntracker. Clock was 26 seconds slow. Reset clock and adjusted Zenith -.3 deg. RTS at 1939 GMT.

8/26/10

JES

NIP and DS are scheduled to be replaced for calibration. R/R *NIP and **DS. Loaded new program at 1357 GMT. RTS @ 1400 GMT.

*Old - 29740E6 *New - 36316E6 **Old - 30812F3 **New - 30811F3

7/29/10

JES

DD is scheduled to be replaced for calibration. R/R *DD. Loaded new program at 1516. RTS @ 1520 GMT. *Old - 34580 *New - 33785

7/22/10

PAD

Suntracker alignment needs to be adjusted. Connected with suntracker. Time is 1 min 48 sec slow. Reset time. Adjusted azimuth -.3 deg. Adjusted Zenith .5 deg. RTS at 1952 GMT.

5/21/10

PAD

BRS DS ventilator fan is not running. R/R ventilator fan at 1845 GMT. RTS at 1935 GMT.

11/3/09

PAD

Flexible conduit is weathered and split between junction box and datalogger box allowing rain to enter conduit. Assisted facilities in R/R conduit at 1934 GMT. Datalogger and DS were briefly disconnected for replacing conduit. RTS at 1949 GMT.

7/6/09

CAW

The DS and NIP need to be replaced for 12 month calibration.

R/R the *DS and **NIP between 1500 GMT and 1520 GMT. Loaded and verified the new program @1522 GMT. RTS @1530 GMT.

*Old - 30946F3 *New - 30812F3 **Old - 30575E6 **New - 29740E6

12/31/08

PAD

Last SDS collection was on 12/30/08 at 1804 GMT. Instruments affected were: mfrsr, mfr10m, and brs. These instruments share a common rocket port. At 1509 GMT, power cycled rocket port located at instruments and verified with computer department instruments would communicate with SDS. RTS at 1524 GMT.

9/19/08

JES

NIP needs to be replaced. R/R *NIP. Loaded new program at 1536 GMT. RTS @ 1540 GMT. *Old - 34504E6 *New - 30575E6

9/12/08

JEM

The *DS, **DD and ***NIP need to be replaced for annual calibration. R/R the radiometers from 1520 GMT to 1535 GMT. Loaded the program @1610 GMT. Verified the SN's and cal factors. RTS @1610 GMT.

*Old - 31636F3 *New - 30946F3 **Old - 33256 **New - 34580 ***Old - 29010E6 ***New - 34504E6

7/29/08

PAD

Found BRS DS radiometer fan had failed during morning PM rounds. R/R fan at 1915 GMT. RTS at 2000 GMT.

3/4/08

PAD

NIP on BRSN is slightly out of alignment with sun. Set Suntracker clock and adjusted Azimuth .2 degrees and Zenith .3 degrees. NIP is now correctly aligned. RTS at 1913 GMT.

11/8/07

MK

Tested the UPS for proper operation. The UPS failed and needs to be replaced. R/R UPS at 1630 GMT.

9/14/07

JEM

PM and monitored data locations. The DS, DD and NIP need to be changed out for annual calibration. R/R the *DD @1540 GMT, **NIP @1535 GMT and ***DS @1535 GMT. Loaded the new program to the datalogger @1543 GMT. Verified all S/N's and Cal Factors. All were correct. RTS @1550 GMT.

*Old - 33238 *New - 33256 **Old - 29851E6 **New - 29010E6 ***Old - 30934F3 ***New - 31636F3

3/1/07

JEM

The suntracker alignment is slightly off. Adjusted the suntracker clock @2044 GMT. The clock was 00:01:10 slow. Also adjusted the zenith by +0.3 and azimuth by +0.1. Alignment is now perfect. RTS @2045 GMT.

8/14/06

PAD

UPS for BRS instruments has been tripping. Removed BRS suntracker from UPS power and plugged into raw A/C power. RTS at 1610 GMT.

7/3/06

CAW

Replaced the *DS, **DD, and ***NIP radiometers for calibration.

*Old - DS 30668F3 *New - 30934F3 **Old - 32972 **New 33238 ***Old - 29850E6 ***New - 29851E6

4/26/06

PAD

At 1456 GMT, found BRS suntracker was not in alignment with sun. Limit switches had stopped movement of instrument. Instrument affected were the DD, DIR, and NIP. At 1500 GMT, Mark Klassen and I manually moved suntracker into position and power cycled instrument. Instrument moved into correct alignment. RTS at 1515 GMT.

4/10/06

MK

The communications equipment needs to be moved and reorganized into a larger enclosure. The UPS and power supplies need to be moved and reorganized into a different enclosure. Disconnected power and communications to all instruments at E13 comms enclosure at 1525 GMT. Moved UPS and power supplies, organized all communications equipment and powered back on at 1945 GMT.

3/21/06

PAD

At 1601 GMT, found BRS radiometers and suntracker without power. Instruments affected were the DS, DD, DIR, and NIP. Power cycled UPS at 1606 GMT. All instruments restarted and suntracker found what appears to be correct alignment. It is overcast and unable to check alignment.

2/16/06

JEM

The upgraded program needs to be loaded onto the Datalogger and Storage Module. Loaded the program to the Datalogger @1903 GMT and to the Storage Module @1904 GMT. RTS @1915 GMT.

12/5/05

MK

The OS needs to be upgraded on all CR10X's. Downloaded current data off of data logger, loaded new OS, retrieved program from data logger, loaded new OS. Loaded program back into CR10X datalogger. System is operating normal. S/N 4648. RTS at 2010 GMT.

9/29/05

MR

The Solar tracker is pointing to the west and not moving. Power cycled the solar tracker at 1521. The tracker rotated around, found the sun and started tracking at 1523 GMT. RTS at 1536.

8/19/05

JEM

The NIP, DD, DS and US need to be changed out for annual calibration. R/R *NIP @1348 GMT, **DD @1350 GMT and ***DS @1352 GMT. Loaded the new DLD to the Datalogger and Storage Module @1415 GMT. Verified all S/N's were correct. RTS @1410 GMT. *Old - 29578E6 *New - 29850E6 **Old - 33278 **New - 32972 ***Old - 30665F3 ***New - 30668F3

8/17/05

MR

The communications between the CF instruments and the SDS has stopped. T/S found that the fiber optic media converter had not reset. Reset the unit in the optical trailer and in the communications box at E13. The communications was restored.

12/17/04

MR

DQPR # 124 requests that the DS on the BRS instrument be checked out due to it consistently reading -10 W\/m^2 at night and 20-40 W/m^2 during the day. Comparing the SIRS instrument wiring with the BRS wiring, I found that one of the grounds was wired to the Ground terminal on the data logger wiring panel instead of the Analog ground on the wiring panel. Moved the ground terminal to the proper location. The DS ventilator fan was also observed as being very dirty and the screen looked plugged. One of the blades was also badly chipped. Upon examining the blades the whole blade assembly sleeve came off with no pressure. R/R the fan and cleaned the screen. This fan looks as if it might be the original one installed in the system. RTS at 2030 GMT.

12/3/04

JEM

The DS, DD and NIP need to be replaced for annual calibration. R/R the *DS @1611 GMT, the **DD @1614 GMT, and the ***NIP @1617 GMT. Verified all serial numbers and coefficients. Loaded the new program to the datalogger and storage module @1655 GMT. RTS @1655 GMT. *Old - 30617F3 *New - 30665F3 **Old - 33247 **New - 33278 ***Old - 31343E6 ***New - 29578E6

9/30/04

JEM

Collected data from the datalogger @1925 GMT due to a collection problem with the SDS. RTS @1938 GMT.

9/7/04

MR

DQPR # 358 requests that the coefficients in the data logger be checked due to unusually high readings from the DIR since 8/27/04 when the SIRS E13 and Sirs test C1 DIR where changed out. T/S found that the wrong data logger programs had been loaded into the wrong instruments. Checked and confirmed which DIR instrument and which program was supposed to be in which data logger. Checked and changed all three instrument (BRS, SIRS E13 and SIRS C1) programs so that the correct coefficients where loaded in the correct location. RTS at 2130 GMT.

8/5/04

MR

The DIR sensor needs to be changed out for calibration. R/R the *DIR. RTS at 1630 GMT. *Old - 30828F3 *New - 30357F3

7/26/04

PAD

At 1400 GMT, sun tracker was found out of position. At 1400 GMT, sun tracker was found out of position. Storms in area on 7/23/04. Power cycled UPS at 1408 GMT. RTS at 1423 GMT.

6/15/04

MR

The NIP instrument is reporting wrong readings possible due to lightning strike observed near the site on 6/2/04. Powered down the instrument at 1955 GMT. R/R the *NIP. Powered up the instrument and downloaded the new program at 2000. RTS at 2010. *Old - 31747E6 *New - 31343E6

6/9/04

CAM

Continued T/S of DQPR 308 (normal incidence values of downwelling shortwave radiation reporting as 9999). Monitored BRS locations and observed NIP reading was most often 9999 and occasionally reported smaller negative values. Disconnected the NIP cable and observed no change in status of the reported data values. Obtained a NIP from the RCF that was due for calibration but known to be good. Connected the NIP to the existing cable and observed that the NIP readings then returned to normal, thus confirming the BRS NIP has failed. Notified Tom Stoffel of the situation and, due to a NIP shortage, he will decide what action to take next. RTS with original NIP in place at 1900 GMT. Note the time of outage corresponds closely with time of a lightning strike on 6/2/04 around 1845 GMT.

6/8/04

CAM

DQPR 308 indicates the normal incidence values of downwelling shortwave radiation are reporting as 9999. Checked all wiring connections at 1800 GMT. No problems found. Will continue T/S tomorrow. Note the time of outage corresponds closely with time of a lightning strike on 6/2/04.

3/16/04

PAD

Instrument mentor requested shorting plug to be installed on DS cable on 3/15/04 for testing. Removed shorting plug from DS cable at 1820 GMT. Attached cable to instrument. RTS at 1835 GMT.

3/16/04

PAD

Instrument mentor requested shorting plug to be installed on DD cable on 3/15/04 for testing. Removed shorting plug from DD cable at 1816 GMT. Attached cable to instrument. RTS at 1831 GMT.

3/15/04

MR

The instrument mentor has requested that the DS and DD signal cables be shorted for testing. Placed shorting plugs on the DS and DD cables. RTS at 2107.

1/8/04

MR

The BRS DS desiccant needs to be changed. Changed the desiccant. RTS at 2145 GMT.

1/8/04

MR

The BRS DIR and DD need to have putty installed on the radiometer cables to keep insects out of the instrument. Installed the putty on the cables. RTS at 2145.

12/11/03

JEM

The DIR needs to be replaced so it can be calibrated. R/R *DIR @1920 GMT. RTS @1935 GMT. *Old - 30687F3 *New - 30828F3

10/29/03

MR

The radiometers are due to be changed out. Shut the system down at 2050 GMT. Changed the *DD, **NIP and ***DS radiometers. Uploaded the new program at 2205 GMT. RTS at 2208 GMT. *Old - 33267 *New 33247 (CF=108.62 W/m2) **Old - 29935E6 **New - 31747E6 (CF=8.34 W/m2) ***Old - 30826F3 ***New - 30617F3 (CF=120.30 W/m2)

9/24/03

MR

DQPR #124 requests that the DS level be checked and the connections on the data logger be reseated. Checked the level and it checked ok. Checked and reseated the connections on the data logger and wiring panel. The connections checked ok. RTS at 1529 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT. Power was restored to the Central Facility at approximately 1945 GMT. RTS @2000 GMT.

12/24/02

JEM

The suntracker has stopped. Power cycled the suntracker @1603 GMT, but the suntracker was unable to move. Powered the tracker off @1604 GMT. A large amount of ice had built up between the NIMFR, suntracker, and the mounting plate for the radiometers that was prohibiting the suntracker from moving. Broke all ice off of the suntracker @1607 GMT. Powered the suntracker back on @1611 GMT. The suntracker appeared to line up in the proper position, but was unable to verify properly alignment due to current sky conditions. RTS @1615 GMT.

11/12/02

MR

The ventilators are not running. T/S found that the dust screens on both ventilators where binding the fans and keeping them from turning. Placed small washer under the corners between the fans and the screens on the mounting holes and this allowed the screens to be raised up off the fans and allowed the fans to run freely. RTS at 1630.

11/7/02

KT

The fan for the DIR ventilator needed to be replaced. At 1530 GMT powered off the suntracker and powered off the BSRN row of instruments. At 1537 GMT removed the DIR ventilator. At 1553 GMT removed and replaced the fan. At 1600 GMT placed DIR back into position. RTS at 1607 GMT. The fan for the DD ventilator needed to be replaced. At 1610 GMT removed the DD ventilator. At 1621 GMT removed and replaced the fan. At 1645 GMT placed DIR back into position. RTS at 1725 GMT.

9/19/02

JEM

The DD, DS, and DIR desiccant needs to be changed. R/R DD, DS, and DIR desiccant. RTS @1648 GMT.

7/3/02

CT

The BSRN DD radiometer ventilator fan was not running. During the inspection of instrument the fuse was found blown. The instrument was powered off at 1915 GMT. The fuse was removed and replaced. Status OK. RTS at 1930 GMT.

7/2/02

CT

The DLD Program needed to be downloaded. The DLD program was downloaded at 2120 GMT. Status OK. RTS at 2135 GMT.

6/27/02

JEM

The *DS, **DD, and ***NIP radiometers need to be changed out for 12 month calibrations. R/R DS radiometer from 1912 GMT to 1918 GMT. R/R DD radiometer from 1922 GMT to 1928 GMT. R/R NIP radiometer from 1919 GMT to 1922 GMT. There is no new program available to load. RTS @1930 GMT. *Old - 31626F3 *New - 30826F3 **Old - 33283 **New 33267 ***Old - 30718E6 ***New - 29965E6

6/17/02

CT

The MFR shadow band was rotating continuously. Power cycled the instrument at 1523 GMT. At 1524 GMT the Bit driver was power cycled. RTS at 1538 GMT. Trouble shooting is in progress.

6/14/02

MR

Power restored. System operational.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT. The instrument is shut down at this time.

5/23/02

KT

The desiccant in the BSRN DS, DIR, and DD radiometers needed to be replaced. Replaced the desiccant in the DS radiometer at 1900 GMT, DIR at 1915 GMT and DD at 1930 GMT. RTS at 1945 GMT.

5/16/02

CT

The BSRN MFR needed to be leveled. The MFR was leveled at 1415 GMT. RTS at 1430 GMT.

5/14/02

MR

The instrument data needs to be collected and placed on the door step for the SDS to collect. Collected the data from 1704-1734 GMT. RTS at 1734 GMT.

5/3/02

CT

The desiccant in the DIR radiometer needed to be replaced. Removed and replaced the desiccant in the DIR radiometer at 1902 GMT. RTS at 1917 GMT.

5/3/02

CT

The desiccant in the DD radiometer needed to be replaced. Removed and replaced the desiccant in the DD radiometer at 1847 GMT. RTS at 1902 GMT.

5/2/02

CT

The BSRN DD radiometer was not level. The BSRN DD radiometer was leveled at 1600 GMT. RTS at 1615 GMT.

4/25/02

CT

The water level on the BSRN MFR battery was low. At 1535 GMT the battery reservoir was filled with water. RTS at 1550 GMT.

4/25/02

CT

The BSRN MFRSR Shadow Band needed to be aligned. Aligned the MFR Shadow Band at 1725 GMT. RTS at 1740 GMT.

4/25/02

CT

The BSRN Suntracker needed to be leveled. The Suntracker was leveled at 1515 GMT. RTS at 1535 GMT.

4/25/02

CT

The Nip alignment needed to be adjusted on the BSRN Suntracker. Adjusted the Zenith by -.4 and the Azimuth by -.5 at 1515 GMT. RTS at 1535 GMT.

3/5/02

CT

The NIP on the BSRN Suntracker was not aligned properly. Aligned the Suntracker NIP by .4 degree zenith and .4 degrees azimuth at 1935 GMT. RTS at 1950 GMT.

2/22/02

CT

The suntracker needed to be realigned. Adjusted the Azimuth on the Suntracker by .5 degrees. RTS at 1609 GMT.

2/12/02

MR

The BRS did not recover from the power outage on 1/30/02. Power cycled the instrument at 1900 GMT. It was observed that the program loaded and restarted at 1903. RTS at 1915 GMT.

2/11/02

MR

The AC power to the site was lost on 1/30/02 due to local ice storm. Powered up the UPS. The BRS instrument reloaded and restarted at 1440 GMT. RTS at 1450 GMT.

12/17/2001

CJ

Desiccant needed to be changed out in the DIR. Removed and replaced desiccant at 1515 GMT. RTS at 1530 GMT.

12/17/2001

CJ

Desiccant needed to be changed out in the DD. Removed and replaced desiccant at 1500 GMT. RTS at 1515 GMT.

12/17/2001

CJ

Desiccant needed to be changed out in the DS. Removed and replaced desiccant at 1445 GMT. RTS at 1500 GMT.

11/20/01

MR

The power to the site will be shut down for approximately 1 hour for a site water leak repair. The data loggers should continue to run due to battery back up but the solar trackers will be down during that time. The trackers powered off at 1500. Upon power being restored to the site the trackers reset itself and returned to the proper orientation. RTS at 1600.

8/15/01

CW

The BSRN radiometers are due for annual calibration change outs. RR *DD, **DS, and ***NIP radiometers. RTS at 1500 GMT. *Old - 33273 *New 33283 **Old 31096F3 **New - 31626F3 ***Old - 30721E6 ***New - 30718E6.

7/18/01

MR

The new multiplexer has come in for the BRS and needs to be installed. Installed the* multiplexer. Powered up the instrument at 1645 GMT. Status ok. RTS at 1700 GMT. *New - 16194

6/26/01

DAN

Removed the *AM-416 from the BRS instrument at 2014 GMT. Jim Liljegren granted approval to remove the mux temporarily to test the Seminole EF SIRS. *Old - 9545

2/27/01

MR

The Card reader keeps corrupting the data card after insertion of the card. Removed the *card reader and will send it back for repairs. *Old - E2507

2/26/01

CAW

The new program needed to be loaded for the new DD radiometer which was installed on 2/22. Loaded the new program and RTS at 1510 GMT.

2/22/01

 JEM

The loaned NREL *DD radiometer needs to be replaced with a CART black and white radiometer. Remove old DD radiometer @ 2037 GMT. Installed new DD radiometer @ 2040 GMT. Did not have the new program to download at this time. The SIRS was functioning properly and returned to service @ 2046 GMT. *Old - 21096 *New - 33273

1/22/01

DN

In accordance with BCR-00288 (BSRN Conversion to BRS or RCT), the following maintenance was performed. Disabled the BSRN collection and ingest processes at 1739 GMT. Secured BSRN solar tracker power at 1843 GMT. BRS solar tracker powered on at 1934 GMT. The BRS instrument was powered on at 2209 GMT following the completion of wiring and testing. BRS collection and ingest processes left disabled pending further instruction. Data will be SneakerNet until the collection and ingest software is ready and a BRS cloud displayed (and functional) on the SGP SDS Clouds Page or other instructions are passed to Site Ops. The *PSP/DD instrument was replaced with an Eppley Black & White instrument. This one is a loaner from NREL and will be replaced with a CART DD when one is available. All other BSRN radiometers remain unchanged for the BRS. The data logger system and programming is essentially identical to that of the SIRS platform, except that inputs for the unused positions of PSP/US and PIR/UIR have been shorted and will be displayed within the data as -9999. New components installed for the conversion include **CR10X, ***AM416 Multiplexer, ****SC32A, *****CSM1 Card Reader, and ******12V Power Supply. *Old - 30953F3 *New - 21096 **New - X4648 ***New - 9545 ****10717 *****E2507 ******New - 12306

8/25/00

KT

At 1440 GMT the desiccant needed to be changed in the DIR, DD, and the DS. R/R desiccant. RTS at 1455 GMT.

8/17/00

KT

The Nip was out of alignment with the sun. At 1403 GMT adjusted the suntracker Azimuth and Zenith. RTS at 1418 GMT. Old Azimuth 91.58, New Azimuth 92.19. Old Zenith 65.30, New Zenith 64.48.

7/28/00

CW

The BSRN radiometers are due for annual calibration change outs. RR *DD, **DS, and ***NIP radiometers. RTS at 1500 GMT. *Old - 30709F3 *New - 30953F3 **Old - 31633F3 **New - 31096F3 ***Old - 30720E6 ***New - 30721E6.

6/15/00

KT

Removed and replaced the desiccant in the DIR, DD, and DS radiometers. RTS at 1459 GMT.

6/2/00

ZW

The BRS DIR ventilator fan was bad. Disconnected AC power to the DIR ventilator fan at 1440 GMT. R/R the vent fan. Reconnected power at 1510 GMT. Status OK. RTS 1510 GMT.

2/11/00

MK

The level, desiccant, ventilators and screens need to be checked on all of the upwelling and downwelling radiometers. Cleaned and checked all BRS radiometers at 1900 GMT.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT the solar tracker shading will be off until power is restored. At 1536 GMT the AC power was restored to all CF instruments. RTS at 1536 GMT.

12/23/99

KT

The desiccant needed to be changed on the DS for the BSRN. At 1622 GMT removed and replaced desiccant. RTS at 1626 GMT.

12/23/99

KT

The desiccant needed to be changed on the DD for the BSRN. At 1615 GMT removed and replaced desiccant. RTS at 1620 GMT.

12/23/99

KT

The desiccant needed to be changed on the DIR for the BSRN. At 1607 GMT removed and replaced desiccant. RTS at 1614 GMT.

11/26/99

KT

At 1925 GMT powered off sun tracker to check ventilator fan for the DIR on the BSR. Noticed a build up of dust and dirt on fan screen. Cleaned fan grill and fan started operating. At 2007 GMT powered sun tracker on. RTS at 2015 GMT.

11/12/99

KT

At 1823 GMT the preamp cable entering into the NIP had 2 exposed wires that needed to be reseated. Reseated the exposed wires and taped so wires would not come loose. RTS at 1828 GMT.

10/26/99

KT

At 1815 adjusted the Azimuth by 1.2 and at 1820 GMT adjusted the Zenith by .04. RTS at 1828 GMT.

10/15/99

MR

RR *DS at mentors request. *Old - 29608 *New - 31636

9/21/99

KT

At 1725 GMT the ventilator shield was found laying on the ground. At 1728 GMT re-installed ventilator shield. RTS at 1731 GMT.

9/8/99

MR

R/R the *data logger due to no files being stored. RTS at 1815. *Old - 27500 *New - 20444.

8/24/99

MR

Reloaded the BSRN program as per instruction of mentor and storage module at 1640 GMT. RTS at 1655.

8/20/99

MK

R/R *DS, **DD, and ***NIP for annual calibration. *Old - 30665 *New - 29608 **Old - 30664 **New - 30709 ***Old - 29743 ***New - 30720

6/23/99

CAM

Leveled DIR at 1605 GMT. Bubble was slightly to the west. Leveled DD at 1608 GMT. Bubble was slightly to the west.

3/25/99

MK

The Suntracker alignment needs to be corrected. Reset the Suntracker time at 2052 GMT. Adjusted the azimuth 0.2 deg. and zenith 0.2 deg. Alignment status normal. RTS at 2100 GMT.

3/2/99

MW

The DD/PSP was out of level. Leveled the DD/PSP 2017-2026 GMT. The DIR/PIR was out of level. Leveled the DIR/PIR 2017-2026 GMT.

2/10/99

ZW

The BSRN Sun Tracker was not level. At 1640 GMT Site Ops leveled the BSRN Sun Tracker. Site Ops also adjusted the azimuth +0.25 degree, and zenith -0.35 degree. Status ok. RTS 1700 GMT.

11/25/98

SS

The desiccant needed changed in the BRS DS, DD, and DIR. Changed the desiccant. RTS at 1720 GMT.

10/22/98

CAM

SDS personnel indicated that 0 byte size files were being received from BSRN since 9/25/98 at 1530 GMT when the radiometers were changed for calibration. In the past this has been an indication that the program has become corrupted and needs to be reloaded. Attempted to monitor the BSRN data logger but all values were zero. Retrieved the program stored in the data logger using PC208W software. Located the SM192 storage module that is kept in spares storage and which contains the original BSRN program. Loaded the program from storage area #8 of the SM192 to a laptop. Compared the program with that which was retrieved from the data logger. The programs were identical with the exception of the data logger EPROM signature. Attempted to store the fresh program to the data logger. The data logger would not accept the program and indicated that the logger was in remote keyboard mode. Installed a new logger and was able to load the program from storage module location #8 into the new logger at 1545 GMT. Brought the original logger to the lab for T/S. Attempted to exit data logger remote keyboard mode by issuing the *0 command. The logger appeared to exit remote keyboard mode and return to telecommunications mode. However each time an attempt was made to load the program it failed and indicated it was still in remote keyboard mode. Meanwhile the BSRN continued to collect 0 byte size files with the new logger as well. Further T/S found that the SM192 storage module actually contains two programs. One is stored in location #8 but there is also another program in location #1. This is different from all other instruments. Reinstalled the original data logger in the BSRN. Connected the SM192 storage module to the data logger and powered the logger on. This forces a dump of the storage module programs (2) into the data logger. Used PC208W software to retrieve the program from the data logger for viewing. The actual program stored in the data logger is a combination of the two programs on the storage module. Apparently the two programs merge together when loaded to the data logger to form a third separate program. RTS status O.K. at 1843 GMT. BSRN collections have been successful since then with normal size files.

9/25/98

ZW

The BSRN DD, DS, and NIP were due to be calibrated. Powered off the BSRN at 1457. R/R the BSRN *DD at 1500 GMT. R/R the BSRN PSP **DS. R/R the BSRN ***NIP. Powered on the BSRN at 1522 GMT Status ok. RTS 1530 GMT. *Old - 30824F3 *New - 30664F3 **Old - 31633F3 **New -30665F3 ***Old - 31387E6 ***New - 29743E6

9/1/98

SS

Daily checks revealed that the desiccant needed changed in the Test Bed and BRS sensors DS, DD, and DIR. Changed the desiccant. RST at 1520 GMT.

8/18/98

MW

The BSRN sun tracker was slightly out of level. Leveled the tracker 1643-1648 GMT.

8/14/98

SS

PM and data collection. PM revealed that the clock was 1:20 minutes slow. Set the clock at 1447 GMT. RTS at 1500 GMT.

8/4/98

MW

The DIR and DD were not level. The tracker was tilting out of level to the Northwest. Adjusted tracker level at 1740 GMT. RTS at 1750 GMT.

8/3/98

ZW

The MFR Mentor came to calibrate the BSRN MFR sensor. At 1815 GMT Site Ops R/R the BSRN MFR sensor head. Powered on system at 1915 GMT. Status ok. RTS 1915 GMT. *Old - 919 *New - 248

6/3/98

MK

The desiccant in the DIR and DD radiometers needs changed out. Changed out DIR and DD desiccant at 1535 GMT. RTS at 1550 GMT.

2/26/98

ZW

The BSRN Suntracker was not aligned perfectly with the Sun. At 2100 GMT using the laptop computer Site Ops aligned the Suntracker perfectly with the Sun by adding .3 degree to the zenith. Status ok. RTS 2120 GMT.

9/17/97

CW/RS

R/R *NIP, **DD, and ***DS with newly calibrated sensors. *Old - 29011 *New - 31387, **Old - 30615 **New - 30824, ***Old - 30617 ***New - 31633.

6/24/97

CW/RS

Changed out *DIR, **DD, ***DS, and ****NIP and test box monitored in each position for 10 minutes. *Old - 28630 *New - 30687, **Old - 30670 **New - 30615, ***Old - 30824 ***New - 30617, ****Old - 29554 ****New - 29011.

05/08/97

MLS

Performed PM. RTS at 1350 GMT.

4/29/97

ZW

The Nip was showing bad readings so removed the Nip cable at 1553 GMT. Made a new cable and replaced it at 1830 GMT. Status ok. RTS 1900 GMT.

4/28/97

ZW

AT 1450 GMT on the daily PM the nip was found out of alignment adjusted Suntracker Azimuth -0.7 and Zenith +0.2. Status ok. RTS 1505 GMT.

6/25/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1346 GMT.

6/14/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1442 GMT.

6/13/96

CAM/GRS

Daily check cleaned dirt from DS radiometer at 1358 GMT.

6/10/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1406 GMT.

6/7/96

CAM

Cleaned dirt from NIP radiometer at 1442 GMT.

6/7/96

MW/ZW

PM. RTS at 1630 GMT.

6/4/96

CAM

Daily check cleaned dirt from NIP radiometer at 1419 GMT.

5/31/96

MR/KT

Daily check cleaned dirt from all downwelling radiometers at 1410 GMT.

5/24/96

KT/ZW

Daily check cleaned dirt from all downwelling radiometers at 1426 GMT.

5/22/96

CAM

Daily check cleaned dirt from NIP radiometer.

5/17/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1440 GMT.

5/10/96

CAM

Daily check cleaned water from all downwelling radiometers at 1510 GMT.

5/9/96

CAM

Daily check cleaned dirt from NIP radiometers at 1415 GMT.

5/6/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1405 GMT.

5/3/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1424 GMT.

4/26/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1403 GMT.

4/22/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1405 GMT.

4/19/96

CAM/MR

Daily check cleaned dirt from all downwelling radiometers at 1451 GMT.

4/17/96

MR

Daily check cleaned dirt from all downwelling radiometers.

4/15/96

CAM

Daily check cleaned dirt from all downwelling radiometers.

4/12/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1403 GMT.

4/9/96

CAM

Daily check cleaned dirt from NIP and DIR radiometers at 1406 GMT.

4/3/96

CAM

Daily check cleaned dirt from all downwelling radiometers at 1643 GMT.

4/1/96

CAM

Daily check cleaned dirt from NIP radiometer at 1503 GMT.

3/28/96

TMG

Daily check cleaned water from NIP radiometer at 1551 GMT.

3/27/96

CAM

Daily check. Cleaned water from all downwelling sensor domes at 1436 GMT.

3/26/96

CAM

Installed a safety cover over the exposed electrical hazard. Removed the safety lock out.

3/25/96

CAM

Daily check. Cleaned all downwelling sensor domes at 1546 GMT.

3/22/96

 

PM.

3/19/96

CAM

Daily check. Cleaned NIP and DS sensor domes at 1506 GMT.

3/18/96

CAM

Daily check. Cleaned NIP and DD sensor domes at 1451 GMT.

3/8/96

ZW

PM.

3/1/96

CAM

Found electrical safety hazard inside the CR10 enclosure box. Applied a safety lock and tag.

2/23/96

ZW

PM.

2/22/96

TMG/CAM

Moved the CR10 enclosure box to new location as part of the upgrade. Re-routed the radiometer sensor data cables through the below ground conduit. RTS at 2148 GMT.

2/13/96

TMG/GRS

AC power secured for construction at 1529 GMT. RTS at 2002 GMT.

2/9/96

 

PM.

2/1/96

TMG/CAM

Daily check. Found solar tracker in "home" position. T/S found nothing. Manually moved the Zenith gear out of the limit switch, power cycled and tested OK. RTS at 1440 GMT.

1/31/96

TMG/CAM

Special visit. T/S coms found EFCOS modem powered off. RTS at 1440 GMT.

 1/29/96

TMG

AC power secured for construction at 1432 GMT. RTS at 2032 GMT.

1/12/96

MW/ZW

PM. RTS at 1840 GMT.

1/8/96

TMG/CAM

Aligned the solar tracker at 1445 GMT.

1/4/96

TMG/CAM

Special visit. Upgraded the BSRN system to conform with the SIROS. *Removed solar tracker computer. **R/R solar tracker and moved the NIP, DIR, and DD sensors. Added ventilators to ***DD and ****DIR. RTS at 2115 GMT. * Old - 401000877 ** Old - 8760 ** New - 940023 *** New - V6-090 **** New - V6-091

12/29/95

MR/ZW

PM.

12/27/95

TMG

AC power secured for construction at 1510 GMT. RTS at 2215 GMT.

12/15/95

MR/KT

PM.

10/13/95

GRS/MW

Special visit. R/R *DS and **DD radiometers and tested OK. RTS at 1955 GMT. * old - 14862F3 * new - 30824F3 ** old - 29617F3 ** new - 30670F3

9/28/95

CAW/CAM

Special visit. T/S coms failure found SC32A bad. R/R and tested OK. RTS at 1630 GMT.

9/21/95

 

PM. * R/R CR10 data logger and tested OK. * old - 09044 * new - 27500

 5/19/95

GRS

PM. RTS at 1935 GMT.